OpenStudy (anonymous):

(tanx+secx)(1-sinx)=cosx

OpenStudy (unklerhaukus):

(tan+sec) * (1-sin) = (sin/cos +1/cos) * (1-sin) = (sin+1)/cos * (1-sin) = (sin+1)(1-sin)/cos =

OpenStudy (unklerhaukus):

simplify the numerator, and use sin^2+cos^2 =1, ie sin^2 = 1-cos^2

OpenStudy (unklerhaukus):

@12357 are you there?

OpenStudy (anonymous):

yes thank you so much! I understand now

OpenStudy (unklerhaukus):

for these trig proof, it is often best to convert all trig function to sines, and cosines first

OpenStudy (anonymous):

ok great thanks! do you know how to solve trig equations such as tan^2x-secx=1 where the answers should be exact

OpenStudy (unklerhaukus):

im not sure about this one, try closing this question and asking that question as a separate post

OpenStudy (anonymous):

ok I will thanks!